OpenStudy (anonymous):

Segment has midpoint P. If M has coordinates (14, -3) and P has coordinates (1, 4), what are the coordinates of N? using the distance formula didnt work. its looking for a set of coordinates

OpenStudy (anonymous):

which segment has midpt. P?

OpenStudy (anonymous):

segment mn the answer given was (-12,11) i have no idea how they got it though.

OpenStudy (anonymous):

xP=(xM+xN)/2 =>xN=2xP-xM=2-14=-12 yP=(yM+yN)/2 =>yN=2yP-yM=8+3=11 P(-12;11)

OpenStudy (anonymous):

thank you